An accounting firm is planning for the next tax preparation season. From last year's returns, the firm collects a systematic random sample of 100 filings. The 100 filings showed an average preparation time of 90 minutes with a standard deviation of 140 minutes. What is the probability that the mean completion time is between 1 and 2 hours, i.e., 60 and 120 minutes?
D. 0.9676

How do I get this number 0.9676

Thanks
Answer #1

Need to find P(60 < x < 120) = ?

z = (60 - 90)/(140/sqrt(100)) = -2.1428
z = (120 - 90)/(140/sqrt(100)) = 2.1428

P(60 < x < 120) = P(-2.1428 < z < 2.1428)
= 0.9678 (Ans.)